The Lin household’s annual income is $188000. Based on the U.S. federal tax rates below, what is the average tax rate for the Lin household? (Assume zero tax deductions for solving this problem.) Choose one: A. 45% B. 24.41% C. 22.22% D. 33%
Answer #1

Solution:

As per the tax bracket for head of household for 2019, which is:

10% for $0 - $13,600

12% for $13,601 - $51,800

22% for $51,801 - $82,500

24% for $82,501 - $157,500

32% for $157,501 - $200,000

35% for $200,001 - $500,000

37% for $500,001 and above

Then, for given Lin's household annual income of $188,000, we can calculate average tax rate as follows:

Taxable income = $188,000

Total taxes = (13600-0)*0.1 + (51800-13600)*0.12 + (82500-51800)*0.22 + (157500-82500)*0.24 + (18800-157500)*0.32

Total taxes = 1360+4584+6754+18000+9760

Total taxes = $40,458

So, the average tax rate = total taxes/taxable income

ATR = 40458/188000 = 0.2152

So, average tax rate is 21.52%. The closest answer to this one is 22.22%, so correct option is C. 22.22%.

(Possible reasons for difference in actual and given answer: I have used federal tax rate for head of household, nothing is mentioned regarding that in the question so tax rate on single or married with joint filing could also work. I have use 2018 tax brackets (latest) but 2017 might have been used for the question. The answer was supposed to be near about to confuse the students. In case any of the mentioned things holds, you could work with the problem accordingly, as the formulation is the same as given in the solution.)